OpenStudy (anonymous):

the distance on the number line from -10 to -2 equals 5x - 12. what is the value of x?

OpenStudy (jackellyn):

-10 - (-2) = 5x-12 -10 + 2 = 5x -12 -8 = 5x -12 What is the next step to get x alone?

OpenStudy (anonymous):

shouldnt it be to add 12 to both sides?

OpenStudy (jackellyn):

exactly. what do you get?

OpenStudy (anonymous):

4?

OpenStudy (jackellyn):

yup, write out the equation that is left over.

OpenStudy (anonymous):

x = 4/5 ?

OpenStudy (jackellyn):

exactly @lolJDlol

OpenStudy (anonymous):

yay!